OpenStudy (anonymous):

How do I do this? Please don't give me the answer. :0 @geerky42 @Kainui @iambatman @SolomonZelman

OpenStudy (anonymous):

The following scatterplot shows the different amounts a babysitter charges, based on the number of hours that she babysits. Which equation best models the babysitter's charges, where x is the time of the babysitting (in hours) and y is the amount that the babysitter charges (in dollars)? A. y = 6x + 10 B. y = 16x C. y = 10x + 6 D. y = 8x

OpenStudy (phi):

I would find the equation of a line between the first and last points. the points appear to lie on a line (close enough)

OpenStudy (phi):

so find the slope between (16,1) and (40,5) then use y - y0 = m(x - x0) where (x0,y0) is any point on the line. example (16,1)

OpenStudy (phi):

There are fancier ways to find a "line of best fit" that minimize the total squared error, but we probably don't need to use those techniques here.

OpenStudy (phi):

remember, slope is change in y divided by change in x

OpenStudy (anonymous):

How do I find the slope between (16,1) and (40,5)

OpenStudy (phi):

oops, I swapped x and y we want to use (1,16) and (5,40)

OpenStudy (phi):

find the change in y. that means take the y value from the second point (the 40) minus the y value from the first point (the 16) change in y is 40 - 16

OpenStudy (anonymous):

24

OpenStudy (phi):

now do the same for the x values to find the change in x

OpenStudy (anonymous):

4

OpenStudy (phi):

are you stuck on finding the change in x= 5-1 ? now find the slope m = change in y divided by change in x

OpenStudy (anonymous):

24/4=6

OpenStudy (phi):

now we use y - y0 = m(x - x0) m is 6 for (x0,y0) we can use (1,16) that means replace x0 with 1 and y0 with 16 in the formula what do you get ?

OpenStudy (anonymous):

X=1 Y=16 A. y = 6x + 10 -----> 16 B. y = 16x ---------> 16 C. y = 10x + 6 -----> 16 D. y = 8x ----------> 8 What number are we looking for? o_o

OpenStudy (phi):

start with y - y0 = m(x - x0) replace x0 with 1 replace y0 with 16 replace m with 6 y - 16 = 6(x - 1) now we simplify that so it looks like one of the choices

OpenStudy (phi):

I would distribute the 6 and change 6(x -1 ) to 6x - 6

OpenStudy (phi):

then add +16 to both sides

OpenStudy (anonymous):

So how would that equation look?

OpenStudy (anonymous):

A or C?

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@marissalovescats @mathmale